The following data refer to two close-coiled helical steel springs.

 

SPRING

FREE HEIGHT

d

D

N

1

55 mm

4,75 mm

30 mm

8

2

48 mm

5,3 mm

44 mm

6

 

 

One spring is placed inside the other, not touching each other, and compressed between a pair of parallel plates until the distance between the plates is 40 mm. Make a neat detailed sketch of the spring and plate assembly. If G = 83 GPa for both springs, calculate the following:

 

5.1 The force applied to the spring assembly. ​​​​​/4/

 

5.2 The shear stress induced in each spring.​​​​​​/3/

 

5.3 The total strain energy stored in the spring assembly.​​​​/3/

 

5.4 The stiffness of each spring.​​​​​​​/3/

 

5.5​The equivalent stiffness of a single spring that can be substituted in place of these springs.​​​​


1
Expert's answer
2021-07-19T03:15:58-0400

"G = 83\ GPa = 83×10^9\ Pa\\
G = 83× 10^3\ N/mm²"


(Spring index)

"C_1" "=D\/d = 30\/4.75=6.3"

"C_2 =D\/d = 44\/5.3 = 8.3"


(Pitch)

"p_1=" free length/N-1 "=55\/8-1 ="

"55\/7=7.9mm"

"p_2= 48\/6-1 = 48\/5 = 9.6mm"


"τ = F×D/2×\pi d² = 83×10³×30/2×π(4.75)² ="

"8.8×10⁴\ Nm"